OpenStudy (driftracer305):

express as single log...:- 4 log x - 6 log (x+2)

OpenStudy (driftracer305):

@hartnn @ganeshie8

OpenStudy (driftracer305):

@Luigi0210 ?

ganeshie8 (ganeshie8):

use below property to start with : \[\large a \log b = \log b^a\]

ganeshie8 (ganeshie8):

apply that for both the terms in your expression, what do u get ?

OpenStudy (driftracer305):

logx^4 and log(x+2)^6

OpenStudy (driftracer305):

wait i knw the property

OpenStudy (driftracer305):

so it is log x ^ 4/(6)......... but wat about base that are different?

OpenStudy (anonymous):

\[\Large \log_{}a-\log_{}b=\log(\frac{a}{b} )\]

ganeshie8 (ganeshie8):

Excellent ! so we have : \[\large 4 \log x - 6 \log (x+2)\] \[ \large = \log x^4 -\log (x+2)^6\]

OpenStudy (driftracer305):

yea........i saw and i typed the property up there.......but different bases so wat next?

OpenStudy (driftracer305):

we cant do log x ^ (4/6)

ganeshie8 (ganeshie8):

actually x and (x+2) are NOT bases here... there is no explicit base given here - you can assume the base is 10

OpenStudy (anonymous):

Think of the x^4 and the (x+2)^6 work as a and b as I've stated before and what @ganeshie8 there's no base given (it's usually just 10 in this case)

ganeshie8 (ganeshie8):

like below : \[\large 4 \log_{10} x - 6 \log_{10} (x+2)\] \[ \large = \log_{10} x^4 -\log_{10} (x+2)^6\]

OpenStudy (driftracer305):

ohh ok........ so lets say..... log 10 ^ (x^4/ (x+2)^6)....right?

ganeshie8 (ganeshie8):

Looks perfect !

OpenStudy (driftracer305):

ok.......it will take me time to expand the (x+2)^6

OpenStudy (anonymous):

I think you can just leave it just like that, does your work say to expand?

ganeshie8 (ganeshie8):

don't even thinka bout it

OpenStudy (aum):

I'd drop the base 10 in the final answer though as it was not in the original problem.

ganeshie8 (ganeshie8):

\[ \large = \log\left( \dfrac{x^4}{ (x+2)^6}\right)\] END of story.

OpenStudy (driftracer305):

k so my options are :- a) 24 log x/x+2 b) log x^4 (x+2)^6 c)log x (x+2)^24 d) none of the above....

OpenStudy (driftracer305):

it is (d).........i think

OpenStudy (anonymous):

Remember \[\large a \log b = \log b^a\]? Yu'll need to use it again

OpenStudy (driftracer305):

@doulikepiecauseidont ........... Yes Sir...

OpenStudy (driftracer305):

@ganeshie8 .........yea so i think it is option (d)........

ganeshie8 (ganeshie8):

i hesitate to tick NOTA usually... could you take a screenshot of question/options and attach if possible ?

OpenStudy (driftracer305):

yes wait.....give me a second....

OpenStudy (driftracer305):

is it viewable????

ganeshie8 (ganeshie8):

ty :) You're right ! it has to be NOTA

OpenStudy (driftracer305):

@ganeshie8 thanks a lot for the help professor....lol......thanks to u too bro @doulikepiecauseidont

OpenStudy (anonymous):

\[\large \log\left( \dfrac{x^4}{ (x+2)^6}\right)=\log\left( \dfrac{x^2}{ (x+2)^3}\right)^{2}=2\log\left( \dfrac{x^2}{ (x+2)^3}\right)\] I was thinking it was A before but it would be this in case you were wondering..

OpenStudy (anonymous):

np (: